If a spring, with a spring constant of 7.4 N/cm, is compressed from an equilibrium length of 28.2 cm to a new length of 24.9 cm, how much potential energy is stored in it?
Answer #1

given things:

k = 7.4 N/cm = 740 N/m

displacement , x = 28.2 - 24.9 = 3.3 cm = 0.033 m

the potential energy of the harmonic oscillator is given by